You start an old record player and notice a bug on the surface close to the edge of the record. The record has a diameter of 12 inches and completes 33 revolutions each minute.
Answer #1

You didn't provided what to calculate. But still I am calculating what can be evaluated using given values.

Revolutions per minute, Frequency, f= 33

We know that, 1rpm=1/60 rps= 1/60 Hz

also, angular velocity, \omega =2 \pi f

thus, \omega =2 \times 3.14 \times 33 \times (1/60) rad/s

\omega =3.454 rad/s

Now given, diameter,d = 12 inches = 12(0.0254) m= 0.3048m

Radius of record, r= d/2= 0.1524 m

Now, Velocity of bug around record, v=r\omega =0.1524 \times 3.454 = 0.5263 m/s

Also, Time period of bud around record, T=1/f

T=1/[33(1/60)]= 60/33= 1.818s